OpenStudy (anonymous):

The tub of a washer goes into its spin-dry cycle, starting from rest and reaching an an- gular speed of 7.5 rev/s in 7.3 s . At this point the person doing the laundry opens the lid, and a safety switch turns off the washer. The tub slows to rest in 11.1 s . During the whole process of starting up and stopping, through how many revolutions does the tub turn? Assume constant angular acceleration while it is starting and stopping. Answer in units of rev

OpenStudy (anonymous):

I've tried and got the wrong answer of 69

OpenStudy (anonymous):

given data variables: \[ \omega_0=0\qquad \omega_1=7.5{\rm rev/s}=7.5\times2\pi\,{\rm rad/s}\qquad t_1=7.3s\\ \omega_2=0\qquad t_2=11.1s\\ \] 1. how many revolutions = what is \(\theta\) from t=0 to t1? \[\omega_1^2=\omega_0^2+2\alpha\theta\\ {\rm where}\quad \alpha=\frac{\omega_1-\omega_0}{t}=?\] plug in and find theta 2. same thing from t=t1 to t2